H_1 = 0 h_{n+1} = (n+1) * h_{n} + n! Find an explicit formula for a generating function of h_n. U...

h_1 = 0

h_{n+1} = (n+1) * h_{n} + n!

Find an explicit formula for a generating function of h_n.

Use the formula to prove that h_{n} = n! * SUM{from k =1 to n} 1/k.
Answer #1

We want to prove that the solution of (A)\quad h_{n+1} =(n+1)\times h_n + n!\,\text{with} \,h_1= 0 is h_n = n!\times\,\sum _{k=1}^{n}\frac 1k

But note that the required solution does not satisfy the given initial condition h_1 = 0. The initial condition should be h_1 = 1.

Here we first write down two or three successive terms, then guess a pattern. Finally we invoke the principle of mathematical induction to assert that our guess is the correct solution of the given problem.

So here initial condition is h_1 = 1, put n=1 in (A)

h_2 = 2\cdot h_1 + 1! = 2\cdot 1 + 1! =2! + 2!/2 = 2!(1+\frac12) =2!\sum _{k=1}^2\frac1k

put n=2 in (A)

h_3 = 3\cdot h_2 + 2! = 3\cdot (2\cdot 1+ 1!) + 2! =3! + 3\times 1! + 2! = 3!(1+\frac12+ \frac 13) =3!\sum _{k=1}^3\frac1k

So this pattern gives a possible solution to the problem. Assume that the pattern is true for n, That is,

h_n = n!\times\,\sum _{k=1}^{n}\frac 1k

Now we shall prove that the pattern is true for n+1

Consider h_{n+1} = (n+1)\cdot h_n + n! = (n+1)\times[n!\times \sum _{k=1}^n\frac1k] +n! = (n+1)\times n!)\sum _{k=1}^n\frac1k + n!

h_{n+1} = (n+1)!\bigg [\sum _{k=1}^n\frac1k+\frac{1}{n+1}\bigg] = (n+1)!\bigg [\sum _{k=1}^{n+1}\frac1k\bigg]

So by the principle of mathematical induction,the result is true for all integral values of n.